OpenStudy (anonymous):

A ball is thrown from a height of 140 feet with an initial downward velocity of 8/fts . The ball's height h (in feet) after t seconds is given by the following. =h−140−8t16t2 How long after the ball is thrown does it hit the ground?

OpenStudy (freckles):

for it to hit the ground the height must be zero

OpenStudy (freckles):

I don't understand your function

OpenStudy (freckles):

but pretend we have the height is h h(t)=5t^2+t and we want to know when h is 0 replace h(t) with 0 0=5t^2+t and solve for t

OpenStudy (anonymous):

t=0,-1/5 @freckles

OpenStudy (freckles):

that is correct for the example function I gave

OpenStudy (ac3):

can't be 0 doesn't make sense

OpenStudy (freckles):

but as I said I do not understand what your function is

OpenStudy (anonymous):

its for a class i didnt write it a teacher did

OpenStudy (ac3):

write it out using the equation thing

OpenStudy (freckles):

she really wrote "=h−140−8t16t2 " because this makes no sense

OpenStudy (anonymous):

h=140-8t-16t^2

OpenStudy (ac3):

ohhhh

OpenStudy (freckles):

ok replace h with 0

OpenStudy (freckles):

since we are given the height is 0

OpenStudy (freckles):

and you are looking for the second 0 not the first 0

OpenStudy (ac3):

then solve for t using factoring or quadratic formula

OpenStudy (anonymous):

A ball is thrown from a height of 140 feet with an initial downward velocity of 8/fts . The ball's height h (in feet) after t seconds is given by the following.

OpenStudy (freckles):

I think your copying and pasting is not working :p

OpenStudy (freckles):

but anyways you already corrected your function for h so we know it now can you replace h with 0 and solve for t?

OpenStudy (anonymous):

ok thnaks =)

OpenStudy (anonymous):

t=2.72 or t=-3.22

OpenStudy (freckles):

those are approximately right and you should be looking for the second zero that zero being t=2.72 (since this occur later)